Consider the reaction:

N2(g) + O2(g) ⇌ 2NO(g)

EQUIL. 0.200 atm 0.100 atm 1.35 x 10^-4 atm

Calculate the equilibrium pressure of NO(g) AFTER the addition of 0.400 atm of O2(g).

A. 7.82 x 10-4 atm

B. 8.43 x 10-4 atm

C. 3.02 x 10-4 atm

D. 5.02 x 10-4 atm

E. 2.23 x 10-4 atm
Answer #1

The equilibrium constant K_P= \dfrac {P^2_{NO}}{P_{N2}P_{O2}}

K_P= \dfrac {\left ( 1.35 \times 10^{-4} \right )^2}{0.200 \times 0.100}

K_P=9.11 \times 10^{-7}

N_2 O_2 NO
Initial pressure (atm) 0.200 0.100 1.35 \times 10^{-4}
Change in pressure (atm) -x 0.400-x 2x
Equilibrium pressure (atm) 0.200-x 0.500-x 1.35 \times 10^{-4}+2x

The equilibrium constant K_P= \dfrac {P^2_{NO}}{P_{N2}P_{O2}}

9.11 \times 10^{-7}= \dfrac {\left ( 1.35 \times 10^{-4}+2x \right )^2}{\left ( 0.200-x \right ) \times \left ( 0.100-x \right )}

9.11 \times 10^{-7}= \dfrac {\left ( 1.35 \times 10^{-4}+2x \right )^2}{\left ( 0.200-x \right ) \times \left ( 0.500-x \right )}

Since the value of Kp is very small, we equate 0.200-x to 0.200 and 0.500-x to 0.500

9.11 \times 10^{-7}= \dfrac {\left ( 1.35 \times 10^{-4}+2x \right )^2} {0.200 \times 0.500}

9.11 \times 10^{-8} =\left ( 1.35 \times 10^{-4}+2x \right )^2

1.35 \times 10^{-4}+2x=0.000301869

x = 8.34 \times 10^{-5}

Hence, the equilibrium pressure of NO(g) AFTER the addition of 0.400 atm of O2(g).

=1.35 \times 10^{-4}+2x = 1.35 \times 10^{-4}+\left ( 2 \times 8.34 \times 10^{-5} \right )=3.02 \times 10^{-4} \: atm

Thus, the option (C) is the answer.